### v1

# GeoPF: Infusing Geometry into Potential Fields for Reactive Planning in Non-trivial Environments

Abstract
Reactive intelligence remains one of the cornerstones of versatile robotics operating in cluttered, dynamic, and human-centred environments. Among reactive approaches, potential fields (PF) continue to be widely adopted due to their simplicity and real-time applicability. However, existing PF methods typically oversimplify environmental representations by relying on isotropic, point- or sphere-based obstacle approximations. In human-centred settings, this simplification results in overly conservative paths, cumbersome tuning, and computational overhead—even breaking real-time requirements. In response, we propose the Geometric Potential Field (GeoPF), a reactive motion-planning framework that explicitly infuses geometric primitives—points, lines, planes, cubes, and cylinders—into real-time planning. By leveraging precise closed-form distance functions, GeoPF significantly reduces computational complexity and parameter tuning effort. Extensive quantitative analyses consistently show GeoPF’s higher success rates, reduced tuning complexity (a single parameter set across experiments), and substantially lower computational costs (up to 2 orders of magnitude) compared to traditional PF methods. Real-world experiments further validate GeoPF’s robustness and practical ease of deployment. GeoPF provides a fresh perspective on reactive planning problems driving geometric-aware temporal motion generation, enabling flexible and low-latency motion planning suitable for modern robotic applications.
Index Terms:
Reactive Planning, Redundant Robots, Manipulation Planning, Motion and Path Planning
I Introduction
Robotics has expanded beyond structured, well-defined factory floors into (i) unstructured, (ii) cluttered, and (iii) dynamic human-centred spaces, as illustrated in Fig. 1. As robots increasingly share workspaces with humans, the need for reactive intelligent navigation solutions becomes paramount. Yet, despite significant advances in motion planning [1], autonomous reactive behavior in settings (i)-(iii) remains an open problem [2, 3, 4]. The core challenge lies in balancing two competing demands: efficiently modeling unstructured and cluttered environments (i)-(ii) while ensuring real-time reactivity to dynamic, unforeseen changes (iii) [5, 6]. This challenge is exemplified in Fig. 1, where a robot must navigate a constrained environment to pour tea, set a breakfast table, and dynamically respond to a human moving objects in real-time. Achieving autonomy and reactive intelligence in such scenarios requires a structured and stable environment representation—one that captures both spatial and temporal information—with a mapping that seamlessly integrates this knowledge into reactive motion generation alongside control constraints and real-time actions. This work addresses this open challenge by providing a novel solution for reactive intelligence, the centerpiece of truly versatile robotic systems operating in human-centered non-trivial environments [7].
Traditional monolithic Sense-Plan-Act (SPA) architectures divide the system into sequential sensing, planning, and action phases [1]. This strong modularization enables robots to function effectively in structured, pre-modeled environments [8, 9, 10]. The key advantage of this paradigm lies in breaking down complex tasks into manageable subproblems, making planning more intuitive and computationally feasible [3]. Most successful approaches including graph-based planners, such as Sparse Roadmaps [11] and Probabilistic Roadmaps (PRMs)[12], construct a free configuration space (C-space) representation and connect edges using local planners, often trading optimality for computational efficiency (e.g., A* [13, 14]). In contrast, tree-based single-query planners, such as Rapidly-exploring Random Trees (RRTs), iteratively expand toward the goal while verifying collision-free paths [15]. Despite their effectiveness in structured settings, these methods lack real-time adaptability. Indeed, while SPA methods excel in handling unstructured (i) and cluttered (ii) scenarios, they fail to adapt efficiently to dynamic (iii) changes [3, 4]. Their inherent sequential nature hinders real-time feedback, making them ill-suited for scenarios where reactive behavior and continuous adaptation to environmental changes are critical [16].
Figure 1: Reactive robot behaviour in unstructured, cluttered, dynamic environments. The robot transfers and pours the tea from the shelf to the cup (left), then returns the set while reacting to a human dynamically moving objects (right)—requiring accurate geometric representations for real-time motion adaptation.
In contrast, vector-field-based motion generation, often termed control-based planning, enables rapid adaptation to environmental changes [17], offering an efficient solution to reactivity (iii). These methods often use gradient-based techniques to guide motion generation [18, 19, 20], ensuring real-time responsiveness. While early implementations suffered from local minima, e.g., artificial potential fields (APFs), these have been mitigated by force-based strategies such as circular fields [21, 22, 23, 4] and magnetic fields [24, 25, 26]. Furthermore, hybrid approaches have attempted to integrate C-space planners with vector-field strategies, leveraging an initial trajectory while benefiting from the reactivity of field-based methods [27, 28]. Learning-based approaches have also been explored to enhance field-based planners [29]. However, all these methods still introduce additional computational complexity without resolving the core bottleneck of field-based methods—their reliance on simplified environment representation.
Despite their computational efficiency compared to SPA, these methods rely on oversimplified workspace representations. Inspired by physics, they often model their interaction via particle-based force dynamics, which assumes point/spherical obstacle representations. While intuitive, this approach poorly represents complex objects like shelves, forcing a trade-off: large repulsive spheres result in overly conservative planning, whereas multiple smaller spheres, ellipsoids, or point clouds lead to excessive distance computations, making real-time execution impractical. Conversely, sparsely selecting sample points causes force discontinuities, complex switching strategies, and short-sighted environment representations, ultimately leading to suboptimal or infeasible trajectories that are highly sensitive to tuning parameters.
In summary, despite the promise of field-based planners in providing intuitive and reactive behaviours, their application to real-world robotic manipulation remains constrained by three fundamental issues: (i) high computational costs associated with representing large, non-isotropic vector fields; (ii) limited performance due to the reliance on isotropic and point-based distance fields; and (iii) the complex and impractical tuning requirements necessary to handle real-world scenarios effectively. These shortcomings are inherently tied to the environment representation upon which existing field-based methods are built. In most cases, obstacles are represented using point clouds [23, 4], or through spheres or octomaps [30, 17, 31] leading to challenges in handling non-spherical objects and non-isotropic vector fields. This results in either excessive computational overhead or complex non-intuitive and cumbersome tuning to handle sparse and changing obstacle representation—both cases leading to poor or even infeasible trajectory generation.
Contributions
In this work, we take a fundamentally different approach to reactive planning by casting geometric primitives for improved environment representation into force generation. Instead of focusing on incremental improvements to field-based methods—refining point-based techniques, improving tuning, or modifying isotropic fields—we propose a geometric-based representation that explicitly incorporates the spatial properties of objects into the temporal motion generation process. Our key contributions include: (C1) A novel geometric-aware environment representation for collision avoidance, leveraging fundamental geometric primitives (points, lines, planes, cubes, and cylinders) to define more structured avoidance strategies; (C2) The design of a reactive planning strategy, the Geometric Potential Field (GeoPF), based on geometric distances, which surpasses existing methods in computational efficiency, trajectory feasibility, and robustness while requiring minimal tuning; (C3) Extensive quantitative comparisons and real-world experiments, demonstrating that GeoPF achieves orders-of-magnitude improvements in computational speed over state-of-the-art reactive field-based planners while maintaining higher success rates and superior trajectory quality.
II GeoPF: System Design and Overview
Figure 2: A snapshot of our GeoPF planner in the same setup as Fig. 1(left). Each obstacle, captured as a sphere, line, plane, cube, or cylinder, generates a distance-based repulsive force. The coloured arrows illustrate the aggregated forces—per geometric type—steering the end-effector (EE) safely to the goal without collision.
This work is about reactive planning in non-trivial human-centric environments. Although straightforward extensions are possible for the whole-body [32], our current investigation focuses on the end-effector (EE) $\underline{\bm{x}}{\in}SE(3)$. Let’s consider an initial and a goal pose, respectively $\bm{\underline{x}}^{0},\bm{\underline{x}}^{g}$. Both rotational and translational components are modelled through unit-dual quaternion parametrization, and while the orientation tracking is achieved through screw interpolation [33, 34], in this work, we are interested in the EE cartesian trajectory critical for collision avoidance. This is modelled as a point mass ${\bm{\underline{p}}^{R}}{\in}\mathbb{R}^{3}$.
We list all additional notations used herein in Table I. Furthermore, to improve the mathematical description, we will explore the following functions.
${\mathcal{U}}(\bm{\underline{p}}){=}\tfrac{\bm{\underline{p}}}{||\bm{\underline{p}}||},\quad{\mathcal{N}}(\bm{\underline{p}}_{1},\bm{\underline{p}}_{2}){=}{\mathcal{U}}(\bm{\underline{p}}^{L}_{1}{-}\bm{\underline{p}}^{L}_{2})=\tfrac{(\bm{\underline{p}}^{L}_{1}{-}\bm{\underline{p}}^{L}_{2})}{||(\bm{\underline{p}}^{L}_{1}{-}\bm{\underline{p}}^{L}_{2})||}.$ (1)
where ${\mathcal{U}}(\bm{\underline{p}})$ and ${\mathcal{N}}(\bm{\underline{p}}_{1},\bm{\underline{p}}_{2})$ are the unit direction along $\bm{\underline{p}}\in\mathbb{R}^{3}$ and from $\bm{\underline{p}}_{2}$ pointing to $\bm{\underline{p}}_{1}$.
TABLE I: Notations used throughout the paper
Notation Description
<>$\cdot$>< ⋅ >, $\times$ Dot (inner) and Cross products
${Geo,S,L,P,B,Y}$ Geometric Primitives (Geo), Sphere (S),
Line (L), Plane (P), Cubic (B), Cylinder (Y)
$\mathcal{T}$ Robot workspace in $\mathbb{R}^{3}$
$\mathcal{T}^{obs}$ Geometric Primitive Set
$\mathcal{T}_{\varnothing}$ Unfeasible Workspace
$\mathcal{T}^{feas}$ Feasible Workspace
$\bm{\underline{a}},\bm{\underline{v}},\bm{\underline{p}}$ Acceleration, velocity, position in Cartesian Space
$\bm{\underline{p}}^{R}$ Robot Position in Cartesian Space
$\bm{\underline{\mu}}$ Unit direction in Cartesian Space
$\bm{\underline{f}}$ Force unit direction
$\bm{\underline{F}}$ Force
$r^{Geo}$ Radius of sphere, circle, cylinder
$\bm{\underline{p}}_{O}^{Geo}$ Center of sphere, circle, cylinder
Overview of the Design Strategy
Given a convoluted obstacle geometric shape, simpler primitives can be fitted to achieve interactive performance. Herein, we make use of—spheres (balls, rounded objects), lines (wires, thin poles), planes (walls, floors), cubes (boxes, crates), and cylinders (pipes, cans)—as these geometric types effectively cover the majority of objects (we elucidate some of them) encountered in real-life scenarios. In the same vein, we designate the space occupied by each type as the spherical space $\mathcal{T}^{S}$, the line space $\mathcal{T}^{L}$, the plane space $\mathcal{T}^{P}$, the cubic space $\mathcal{T}^{B}$, and the cylinder space $\mathcal{T}^{Y}$. $\mathcal{T}_{obs}$ can now be explicitly defined as $\mathcal{T}^{S}\cup\mathcal{T}^{L}\cup\mathcal{T}^{P}\cup\mathcal{T}^{B}\cup\mathcal{T}^{Y}$. Thus, considering the full workspace of the robot $\mathcal{T}\subset SE(3)$, the feasible space $\mathcal{T}^{feas}$ in a non-trivial environment is $\mathcal{T}^{feas}=\mathcal{T}\setminus(\mathcal{T}_{obs}\cup\mathcal{T}_{\varnothing})$, where $\mathcal{T}_{\varnothing}$ designates the infeasible workspace, e.g., singular-poses, self-collision, joint limits, and workspace out-of-boundaries.
Each geometric primitive is defined by a set of points that describe its spatial configuration in $SE(3)$. Given the geometric shape and pose information, a geometric model can be constructed for each obstacle in the environment. One key feature to this aim is the normal vector, denoted as $\bm{\underline{\mu}}^{Geo}$, where $Geo\in\{S,L,P,B,Y\}$ represents the specific geometric type.
For each geometric primitive, exploring $\bm{\underline{\mu}}^{Geo}$, we are able to project the robot onto the obstacles. The projected point is the perpendicular foot $\bm{\underline{p}}^{Geo}_{\bot R}$. Such a foot system indicates the spatial relationship between them, which in turn defines the avoidance strategy, as depicted in Figs. 3-7.
The core principle of potential fields is to model the robot and environment with attractive and repulsive potentials. The attractive force is along the direction from robot pointing to goal position $\bm{\underline{f}}^{attr}={\mathcal{N}}(\bm{\underline{p}}_{g},{\bm{\underline{p}}^{R}}).$ The magnitude of the force is controlled by a tuning parameter $k^{attr}$. The attractive force exerted on the robot is given by $\bm{\underline{F}}^{attr}=k^{attr}\bm{\underline{f}}^{attr}.$ Differently, the repulsive force $\bm{\underline{f}}^{Geo}$ respects geometric considerations and it is directed along the shortest path from the obstacle to the robot. The magnitude of the repulsive force depends on the length of shortest distance $d^{Geo}$ with a tuning parameter $k^{Geo}$:
$\bm{\underline{F}}^{Geo}=\frac{k^{Geo}}{d^{Geo}}\bm{\underline{f}}^{Geo}.$ (2)
Therefore, the force is directly proportional to the obstacle distance. The repulsive force should be only active when the robot is close to the obstacle. We set a threshold ${\Delta}^{Geo}$ for the activation. $\bm{\underline{F}}^{\varnothing}$ stems from additional workspace limit-based forces such as self-collision, joint limits, singularities, and workspace limits. The limit-based forces can be expressed through geometric primitives, $\mathcal{T}^{\varnothing}\subseteq\mathcal{T}^{Geo}$. An example is shown in Fig. 11, where boundary limits restrict the feasible workspace.
Problem Statement The robot motion planning problem that we consider is hinged on the notion of potential functions. After the field construction we present earlier, the resultant net force acting on the robot can be expressed as
$\bm{\underline{F}}^{res}=\bm{\underline{F}}^{attr}+\sum_{Geo}\bm{\underline{F}}^{Geo}+\sum_{\varnothing}\bm{\underline{F}}^{\varnothing},$ (3)
The robot’s acceleration at each timestep can be computed as $\bm{\underline{a}}_{i}=\bm{\underline{F}}^{res}_{i}/{m}$, where ${m}$ denotes the mass of the robot, and $i\in\{1,2,3,...\}$ is the timestep index. The velocity will be updated through $\bm{\underline{v}}_{i+1}=\bm{\underline{v}}_{i}+{\delta_{t}}\bm{\underline{a}}_{i}$, ${\delta_{t}}$ being the step size. Thus, the Cartesian robot position updates as
$\bm{\underline{p}}^{R}_{i+1}=\bm{\underline{p}}^{R}_{i}+\frac{1}{2}{\delta_{t}}^{2}\bm{\underline{a}}_{i}+{\delta_{t}}\bm{\underline{v}}_{i}.$ (4)
Lastly, the orientation of the robot is determined by the task requirements, such as the grasping direction. Thus, we seek a feasible trajectory $\mathbb{P}$ starting from $\bm{\underline{x}}^{0}$ and ending in $\bm{\underline{x}}^{g}$ in the free task space $\mathcal{T}^{feas}$ brought about by $\bm{\underline{F}}^{res}$ and taking into account the workspace (environmental) constraints as well as geometrical considerations.
III From Geometry to Force Generation
One of the key contributions of this work is a geometric-based environment representation that explicitly encodes the spatial properties of objects and uses this geometry to drive temporal motion generation. Rather than relying solely on point clouds or large sets of spheres for obstacle modeling, we cast geometric primitives—including spheres, lines, planes, cubes, and cylinders—into repulsive forces within a vector field.
Formally, let $\{G_{i}\}$ denote the set of geometric primitives representing the obstacles. Each primitive $G_{i}$ is associated with its own distance function $d_{i}(\cdot)$ and corresponding repulsive force $\bm{\underline{F}}_{i}$. The aggregate motion is determined by summing these repulsive forces with an attractive force pulling the robot toward the goal $\bm{\underline{p}}^{g}$. Specifically,
$\bm{\underline{F}}^{\mathrm{res}}=\underbrace{k^{attr}\bm{\underline{f}}^{attr}}_{\text{goal attraction}}+\sum_{i}\negthickspace\negthickspace\negthickspace\negthickspace\underbrace{\left(\frac{k_{i}}{\,d_{i}(\cdot)\,}\bm{\underline{f}}_{i}\right)}_{\text{repulsion from obstacle }i}\negthickspace\negthickspace\negthickspace+\ \sum_{j}\bm{\underline{F}}^{\varnothing}_{j},$ (5)
where $k_{\mathrm{attr}}$ and $k_{i}$ are scalar gains. The following subsections detail how each geometric primitive defines its distance function $d_{i}(\cdot)$ and force direction $\mathbf{f}_{i}$.
III-A Point and Sphere-based Potential Field
A sphere is defined by its centroid ${\bm{\underline{p}}^{S}_{O}}{\in}\mathbb{R}^{3}$ and radius ${r^{S}}$, while a point is given by ${r^{S}}{=}0$. It forms the simplest geometric primitive we define for collision avoidance. First, we compute the shortest distance ($d^{S}$) between the point robot and the sphere’s surface, i.e., $d^{S}{=}||\bm{\underline{p}}^{R}-{\bm{\underline{p}}^{S}_{O}}||-{r^{S}}$. Second, we are interested in the direction of the relative position vector, i.e., the radial vector outward from the sphere center, $\bm{\underline{f}}^{S}{=}{\mathcal{N}}(\bm{\underline{p}}^{R},{\bm{\underline{p}}^{S}_{O}})$. The repulsive force is therefore computed as
$\bm{\underline{F}}^{S}=\tfrac{k^{S}}{d^{S}}\bm{\underline{f}}^{S},\quad\text{where }\bm{\underline{f}}^{S}{=}\tfrac{\bm{\underline{p}}^{R}{-}{\bm{\underline{p}}^{S}_{O}}}{||\bm{\underline{p}}^{R}{-}{\bm{\underline{p}}^{S}_{O}}||},\ \ k^{S}{>}0.$ (6)
The sphere-based repulsive force can be enhanced by circular forces or magnetic-based [26, 4], yet we explicitly decided to maintain a straightforward approach, avoiding unnecessary complexity of the spherization—that often rely on tuning for efficiency, as shown in Section IV-A. For more details on alternative forces, please refer to [4, 23].
III-B Segment Line-based Potential Field
In the remainder of this section, we present additional geometric-primitive-based potential field generation. Starting from a line segment, defined by two vertices $\bm{\underline{p}}^{L}_{1},\bm{\underline{p}}^{L}_{2}{\in}\mathbb{R}^{3}$.
The repulsive force surrounding the line segment is a function of the shortest path from the line to the robot. The shortest distance, and thereafter, the repulsive force direction, between the robot ($\bm{\underline{p}}^{R}$) and the finite segment may arise from one of two cases, as shown in Fig. 3.
  1. 1.
    The robot projected onto the line segment, as shown in Fig 3(a), indicating that the perpendicular foot lies between the vertices. In the other words, the robot is orthogonal to the line segment.
  2. 2.
    Otherwise, the robot lies outside the orthogonal region, as shown in Fig 3(b). In the other words, the robot is on the side of the line segment.
In the first case, the resulting orthogonal-based force direction is defined as $\bm{\underline{f}}^{L}_{o}$, whereas the side forces are given by $\bm{\underline{f}}^{L}_{s1}$ and $\bm{\underline{f}}^{L}_{s2}$, as shown in Fig. 3.
To evaluate both conditions, we first compute the perpendicular foot that the robot projects onto the line. To compute it, we first obtain the robot-line plane based on the line within the plane, e.g., $\bm{\underline{\mu}}^{L}_{1\to R}{=}{\mathcal{N}}(\bm{\underline{p}}^{R},\bm{\underline{p}}^{L}_{1})$, and the plane’s normal direction, given by $\bm{\underline{\mu}}^{RL}{=}(\bm{\underline{\mu}}^{L}_{1\to R}{\times}\bm{\underline{\mu}}^{L})$, where $\bm{\underline{\mu}}^{L}{=}\bm{\underline{\mu}}^{L}_{2\to 1}{=}{\mathcal{N}}(\bm{\underline{p}}^{L}_{1},\bm{\underline{p}}^{L}_{2})$ is the unitary vector from $\bm{\underline{p}}^{L}_{2}$ towards $\bm{\underline{p}}^{L}_{1}$. Next, we compute the orthogonal direction perpendicular to both the line and the plane normal, i.e., $\bm{\underline{\mu}}^{L}_{o}=(\bm{\underline{\mu}}^{L}{\times}\bm{\underline{\mu}}^{L}_{R})$. Finally, the orthogonal distance from the robot to the line is given by $d^{L}_{o}=<\bm{\underline{p}}^{R}{-}\bm{\underline{p}}^{L}_{1},\bm{\underline{\mu}}^{L}_{o}>$, which is the length of the projection of the robot to $\bm{\underline{p}}^{L}_{1}$ over $\bm{\underline{\mu}}^{L}_{o}$. The perpendicular foot, in this case, also the projection point, can be calculated as $\bm{\underline{p}}_{\bot R}^{L}=(\bm{\underline{p}}^{R}{-}d^{L}_{o})\bm{\underline{\mu}}^{L}_{o}$.
When the perpendicular foot $(\bm{\underline{p}}_{\bot R}^{L})$ lies between the vertices of the line segment, i.e., $\bm{\underline{p}}_{\bot R}^{L}{\in}\mathcal{T}^{L}$, the repulsive force is the orthogonal force. In this case, the distance from the perpendicular foot to the edge is always smaller than the line distance, i.e., $d^{L}_{1}<l^{L}\cap d^{L}_{2}>l^{L}$. Note that the distance can be calculated using the Euclidean norm, given the known position of perpendicular foot and the vertices. Conversely, when the perpendicular foot is outside the line segment $\bm{\underline{p}}_{\bot R}^{L}\not\in\mathcal{T}^{L}$, the repulsive force is the side force, directed from the closest vertices towards the robot:
[…]
$\bm{\underline{F}}^{L}=\tfrac{k^{L}}{d^{L}}\bm{\underline{f}}^{L}.$ (7d)
((a)) Orthogonal Line Field
((b)) Side Line Field
Figure 3: Line-based potential field. (a) When the perpendicular foot is inside the line segment. The force is orthogonal. (b) When the perpendicular foot is outside the segment, the force is side force.
III-C Plane-based Potential Field
Planes are among the most common geometric obstacles in human-centred environments—representing tables, walls, and other flat surfaces. Traditional methods approximate them using multiple discrete spheres, as shown in Fig. 8(b). Although conceptually simple, this results in inefficient and imprecise representations, requiring extensive tuning to prevent the field from either penetrating the spheres or generating unnecessarily large repulsive forces. In contrast, explicitly considering a plane with four vertices allows for closed-form distance and force computations, reducing both computational overhead and tuning complexity. Indeed, let’s consider a plane defined by four (counter-)clockwise ordered vertices $\bm{\underline{p}}^{P}_{1},\bm{\underline{p}}^{P}_{2},\bm{\underline{p}}^{P}_{3},\bm{\underline{p}}^{P}_{4}{\in}\mathbb{R}^{3}$. The plane’s outward normal can be computed as
$\bm{\underline{\mu}}^{{P}}={\mathcal{U}}(\bm{\underline{\mu}}_{2\to 1}^{P}\times\bm{\underline{\mu}}_{2\to 3}^{P}),$ (8)
where $\bm{\underline{\mu}}_{2\to 1}^{P}={\mathcal{N}}(\bm{\underline{p}}^{P}_{1},\bm{\underline{p}}^{P}_{2})$ and $\bm{\underline{\mu}}_{2\to 3}^{P}={\mathcal{N}}(\bm{\underline{p}}^{L}_{3},\bm{\underline{p}}^{L}_{2})$ depict segments of line within the plane.
Similar to the segment-line case, the repulsive force, thus the avoidance strategy, is determined by the robot’s projection onto the perpendicular foot ($\bm{\underline{p}}_{\bot R}^{P}$) on the obstacle plane.
The perpendicular foot on the plane is computed by $\bm{\underline{p}}^{P}_{\bot R}=\bm{\underline{p}}^{R}{-}d^{P}_{o}\bm{\underline{\mu}}^{P}$ where $d^{P}_{o}={<{\bm{\underline{\mu}}^{P},\bm{\underline{p}}^{R}{-}\bm{\underline{p}}^{P}_{1}}>}$ is the spatial distance between the plane and the robot projected on the orthogonal outward normal from (8).
((a)) Orthogonal Plane Field
((b)) Side Plane Field
Figure 4: Plane-based potential field defined by the perpendicular foot $\bm{\underline{p}}^{P}_{\bot R}$. (a) If it lies within the plane boundary, the condition vectors (green arrows) remain aligned, see (10), then the repulsive force direction $\bm{\underline{f}}^{P}$ aligns with the plane’s normal. (b) Otherwise, the force is derived from the nearest edge.
Whether the perpendicular foot lies within a rectangle’s plane can be evaluated by angles between the lines connecting the foot point and the rectangle’s vertices. Let the direction from the perpendicular foot to the $i^{th}$ vertex be denoted as $\bm{\underline{\mu}}^{P}_{ri}$, where $i\in\{1,2,3,4\}$. If the perpendicular foot is located within the plane, the angles between two adjacent foot-vertex lines should be less than 180°. In contrast, if any of the angles exceed 180°, the point lies outside the plane.
The angle condition can be determined through the cross-product $<\negthinspace\bm{\underline{\mu}}^{P}_{ri},\bm{\underline{\mu}}^{P}_{r(i+1)}\negmedspace>$. Therefore, we have the normal direction of the plane given by
$\bm{\underline{c}}_{1}:={\mathcal{U}}(\bm{\underline{\mu}}^{P}_{r1}\times\bm{\underline{\mu}}^{P}_{r2}),\quad\bm{\underline{c}}_{2}:={\mathcal{U}}(\bm{\underline{\mu}}^{P}_{r2}\times\bm{\underline{\mu}}^{P}_{r3}),$ (9)
$\bm{\underline{c}}_{3}:={\mathcal{U}}(\bm{\underline{\mu}}^{P}_{r3}\times\bm{\underline{\mu}}^{P}_{r4}),\quad\bm{\underline{c}}_{4}:={\mathcal{U}}(\bm{\underline{\mu}}^{P}_{r4}\times\bm{\underline{\mu}}^{P}_{r1}).$ (10)
The perpendicular foot lies inside the plane if and only if all indicators $\bm{\underline{c}}_{i}$ point in the same direction, i.e., $\bm{\underline{c}}_{1}{=}\bm{\underline{c}}_{2}{=}\bm{\underline{c}}_{3}{=}\bm{\underline{c}}_{4}$, as seen in Fig. 4(a). In this case, the perpendicular foot lies inside the plane boundaries $\bm{\underline{p}}^{P}_{\bot R}\in\mathcal{T}^{P}$, and the repulsive force direction is orthogonal to the plane and aligned to (8), that is, $\bm{\underline{f}}_{o}^{P}=\bm{\underline{\mu}}^{P}=\bm{\underline{c}}_{i}$.
When any of $\bm{\underline{c}}_{i}$ points in the opposite direction (Fig. 4(b)), the point is outside the plane, $\bm{\underline{p}}^{P}_{\bot R}\not\in\mathcal{T}^{P}$, which guarantees the convexity of the plane. The resulting side force $\bm{\underline{f}}_{s}^{P}$ and corresponding distance $d_{s}^{P}$ is then redirected from the nearest plane edge or vertex using line-based logic as in Sec. III-B.
The resulting plane repulsion is computed as
$\bm{\underline{F}}^{P}=\tfrac{k^{P}}{\,d^{P}}\bm{\underline{f}}^{P},$ (11)
where $(\bm{\underline{f}}^{P},d^{P})$ is given by either $(\bm{\underline{f}}_{o}^{P},d_{o}^{P})$ or $(\bm{\underline{f}}_{s}^{P},d_{s}^{P})$ according to the case described above.
Plane-based Correction Potential Field Heuristic
While successful in avoiding collision with plane-based geometries, when a large plane is positioned between the robot and the goal, the attractive force counteracts the repulsive force, causing a trap condition (Fig. 5(a)).
If the line from robot to goal intersects the plane interior, we project a repulsive force parallel to the plane’s surface (toward the nearest plane edge) rather than purely along the normal. This prevents the attractive force from canceling out the plane’s repulsion, allowing the robot to “slide” around or above the plane.
Given the unit vector along the robot-goal line, $\bm{\underline{\mu}}^{RG}={\mathcal{U}}(\bm{\underline{p}}^{R}-\bm{\underline{p}}^{G})$, the intersection point in the unbounded plane is given by $\bm{\underline{p}}^{P}_{inter}{=}\bm{\underline{p}}^{R}{-}d^{P}_{inter}\bm{\underline{\mu}}^{RG}$, where $d^{P}_{inter}{:=}\frac{d^{P}_{o}}{<\bm{\underline{\mu}}^{RG},\bm{\underline{\mu}}^{P}_{o}>}$ is the distance from the robot to the intersection point. If the intersection point lies within the plane, we project a force parallel to the plane toward the nearest plane edge.
By defining the nearest edge direction $\bm{\underline{\mu}}^{P}_{e}$, the parallel direction is given by $\bm{\underline{\mu}}^{P}_{corr}=\bm{\underline{\mu}}^{P}\times\bm{\underline{\mu}}^{P}_{e}$. To avoid crossing toward the farthest edge, see Fig. 5(b), we evaluate the direction from the robot to the nearest edge, denoted as $\bm{\underline{\mu}}^{PR}_{e}$, which corresponds to the opposite direction of the line-to-robot vector, Section III-B. By computing $s=\text{sign}(<\bm{\underline{\mu}}^{P}_{corr},\bm{\underline{\mu}}^{PR}_{e}>)$, we can determine whether the parallel direction toward the nearest edge or its opposite. When $s=1$, the direction is correctly toward the nearest edge ($\alpha<\frac{\pi}{2}$), whereas $s=-1$ indicates the farthest edge ($\alpha>\frac{\pi}{2}$), see Fig. 5. Accordingly, the corrected force is
$\bm{\underline{f}}^{P}_{corr}=s\bm{\underline{\mu}}^{P}_{corr},\quad\text{with}\ \ \bm{\underline{\mu}}^{P}_{corr}{=}\bm{\underline{\mu}}^{P}{\times}\bm{\underline{\mu}}^{P}_{e},$ (12)
with the magnitude of the force according to the distance between the robot and the plane - the orthogonal force and the side force corresponding distance.
((a)) Fail to avoid the plane
((b)) Plane Field Correction
Figure 5: Plane force correction. (a) Without correction, a large plane may trap the robot by canceling the repulsive and attractive forces. (b) When the line connecting the robot and goal intersects with the plane ($\bm{\underline{p}}_{inter}^{P}$), the force is redirected to the parallel direction $\bm{\underline{\mu}}^{P}_{corr}$. If the parallel direction points aways of the nearest edge, the force is redirected according to $\alpha$ with $\bm{\underline{f}}^{P}_{corr}$ being the corrected direction.
orders-of-magnitude improvements in computational speed over state-of-the-art reactive field-based planners while maintaining higher success rates and superior trajectory quality.orders-of-magnitude improvements in computational speed over state-of-the-art reactive field-based planners while maintaining higher success rates and superior trajectory quality.
III-D Geometric Cube-Based Potential Field
Extending from previous primitives, many everyday objects, from boxes to shelves and cabinets, can be approximated as cubes or rectangular prisms. To model a cube, we use eight (counter-)clockwise ordered vertices $\bm{\underline{p}}^{B}_{i}{\in}\mathbb{R}^{3}$,  i={1,…,8}. The vertices $\bm{\underline{p}}^{B}_{1,2,3,4}$ represent one surface, whereas $\bm{\underline{p}}^{B}_{5,6,7,8}$ representing the opposite surface. Moreover, $\bm{\underline{p}}^{B}_{1}$ is aligned with $\bm{\underline{p}}^{B}_{5}$, as illustrated in Fig. 6. Therefore, the cube can be decoupled into six different surfaces $\mathcal{T}^{B}_{S1}$ to $\mathcal{T}^{B}_{S6}$ and their intersection edge $\mathcal{T}^{B}_{Si}\cap\mathcal{T}^{B}_{Sj}$, with $i\neq j\in\{1,2,3,4,5,6\}$.
In practice, each of the six faces is treated like an individual plane (as in Section III-C), so the force generation relies on choosing whichever face is nearest, as shown in Fig. 6(a). When the robot is closest to a cube edge $\mathcal{T}^{B}_{Si}{\cap}\mathcal{T}^{B}_{Sj}$, then the forces generated by $\mathcal{T}^{B}_{Si,j}$ are equivalent and the distance and direction of repulsion are handled via plane or line logic, whichever is more appropriate according to Section III-B. In summary,
$\bm{\underline{F}}^{B}=\frac{k^{B}}{d^{B}}\bm{\underline{f}}^{B},(\bm{\underline{f}}^{B},d^{B})=(\bm{\underline{f}}^{B}_{Si},d^{B}_{Si}),i=\arg\min^{6}_{i=1}d^{B}_{Si}.$ (13)
((a)) Cubic Surface Field on $\mathcal{T}^{B}_{S1}$
((b)) Cubic Edge Field on $\mathcal{T}^{B}_{S2}\cap\mathcal{T}^{B}_{S3}$
Figure 6: Cubic force relies on the closest surface. (a) When robot closest to surface, the surface plane provides the force (Sec. III-C. (b) When robot closest to an edge, either surface can be chosen to calculate the cubic force.
III-E Cylinder-Based Potential Field
Cylinders commonly arise from objects such as cups, bottles, cans, pipes, or cylindrical machinery components. We define a cylinder by its central axis (stemming from two vertices $\bm{\underline{p}}^{Y}_{O1},\bm{\underline{p}}^{Y}_{O2}{\in}\mathbb{R}^{3}$) and the radius $r^{Y}$. We explore the curved surface surrounding the cylinder, and two circular-cross section planes centred at the vertices. The normal direction of the curved surface is the radial direction from the central axis extending outward, Fig 7. The normal directions of the circular surfaces are perpendicular to the plane and point outside the cylinder, then we directly apply the force along to the axis direction $\bm{\underline{f}}^{Y}_{X}$.
When the robot is “beside” the curved surface—as in Fig. 6(a)—the force can be computed akin to a line-based approach, whereas if it is above or below, we treat the top or bottom circle similar to the plane-like surface.
Assuming no penetration, the former occurs when the distance ($d^{Y}_{\bot X}$) from the robot to its perpendicular foot within the cylinder-axis is larger than the radius, whereas $d^{Y}_{\bot X}{<}r^{Y}$ indicates the robot is right above or below the cylinder (Fig. 7(b)).
To compute the resulting forces, we compute a robot-axis plane and the orthogonal vector to the cylinder axis ($\bm{\underline{\mu}}^{Y}_{\bot X}$)—as in Section III-B. Note $d^{Y}_{\bot X}$ is computed along $\bm{\underline{\mu}}^{Y}_{\bot X}$.
Surface-Line Potential Field
When the robot is “beside” the curved surface ($d^{Y}_{\bot X}>r^{Y}$), it is closest to a line segment on the surface of the cylinder. To this aim, we shift the cylinder axis to the closest point of the surface by translating the vertices along the direction $\bm{\underline{\mu}}^{Y}_{\bot X}$, i.e., $\bm{\underline{p}}^{Y}_{Si}{=}\bm{\underline{p}}^{Y}_{Oi}+d^{Y}\bm{\underline{\mu}}^{Y}_{\bot X}$, $\ i{=}\{1,2\}$.
The resulting force is defined by the robot to the segment line $\overline{\bm{\underline{p}}^{Y}_{S1}\bm{\underline{p}}^{Y}_{S2}}$ as in Section III-B. Thus, we have the surface-line orthogonal force $\bm{\underline{f}}^{Y}_{Xo}$, side force $\bm{\underline{f}}^{Y}_{Xs}$ and the surface-line perpendicular foot $\bm{\underline{p}}^{Y}_{\bot X}$.
When the perpendicular foot of surface line $\bm{\underline{p}}^{Y}_{\bot X}$ is in between $\bm{\underline{p}}^{Y}_{S1}$ and $\bm{\underline{p}}^{Y}_{S2}$, the cylinder force is the orthogonal force $\bm{\underline{f}}^{Y}=\bm{\underline{f}}^{Y}_{Xo}$, which is the same as the axis orthogonal vector $\bm{\underline{f}}^{Y}_{Xo}=\bm{\underline{\mu}}^{Y}_{\bot X}$. When the surface line foot $\bm{\underline{p}}^{Y}_{\bot X}$ is outside, the cylinder force is the side force $\bm{\underline{f}}^{Y}=\bm{\underline{f}}^{Y}_{Xsi}$, with $i=\{1,2\}$. The distance is computed from the robot to surface’s line.
((a)) Surface Line Field
((b)) Circular Plane Field
Figure 7: Cylinder Force has two different situations. (a) When the robot is close to the curved surface, the force can be regarded as a line force. (b) When the robot is close to the circular plane, the force can be regarded as the plane force. By checking the circular edge points, we can build the plane field on the circular plane.
Circular-Plane Potential Field
When robot-axis distance is smaller than the radius ($d^{Y}_{\bot X}<r^{Y}$), the robot remains either on the top or bottom of the cylinder (Fig. 7(b)). When the circular plane perpendicular foot is inside the plane: $\bm{\underline{p}}^{Y}_{\bot C}\in\mathcal{T}^{Y}_{C}:||\bm{\underline{p}}^{Y}_{\bot C}-\bm{\underline{p}}^{Y}_{O1}||<r^{Y}$, we apply the circular plane force. Otherwise, it is the surface line force. The circular plane normal is the unit vector along cylinder axis $\bm{\underline{f}}^{Y}_{X}={\mathcal{N}}(\bm{\underline{p}}^{Y}_{O1},\bm{\underline{p}}^{Y}_{O2})$. Note that we have surface points $\bm{\underline{p}}^{Y}_{S1,2}$. Suppose the robot is closer to the upper surface with the surface point $\bm{\underline{p}}^{Y}_{S1}$. Then it is easy to find another point on the circular plane by taking the cross product $\bm{\underline{\mu}}^{Y}_{add}=<\bm{\underline{\mu}}^{Y}_{\bot X},\bm{\underline{f}}^{Y}_{Xo}>$. Then, the four points on the edge of the circular plane are $\bm{\underline{p}}^{Y}_{E1,3}=\bm{\underline{p}}^{Y}_{O1}\pm d^{Y}\bm{\underline{\mu}}^{Y}_{add}$, and $\bm{\underline{p}}^{Y}_{E2,4}=\bm{\underline{p}}^{Y}_{O1}\pm d^{Y}\bm{\underline{\mu}}^{Y}_{\bot X}$. With the three of them, we can shape a plane and calculate the plane force according to Sec. III-C. However, a key distinction is that the angle-based condition for verifying the intersection point $\bm{\underline{p}}^{Y}_{C_{\times}}$ is no longer applicable, as the original plane is circular. Instead, the intersection point can be checked by the distance. If the distance from the intersection point to the circle center exceeds the radius $\bm{\underline{p}}^{Y}_{C_{\times}}\notin\mathcal{T}^{Y}_{C}:||\bm{\underline{p}}^{Y}_{C_{\times}}-\bm{\underline{p}}^{Y}_{O1}||>r^{Y}$, the intersection point lies outside the circular plane. Here, the distance is same as the plane-based distance for either correction or non-correction force.
IV Quantitative Analysis and Experiments
In this section, we present an extensive evaluation of our proposed GeoPF in both simulated and real-world settings. Our goal is twofold: (i) To quantitatively assess performance [^1] Key performance metrics include (i) success rate—whether the robot reaches the goal without collision; (ii) computational time—the total time (or average per-step time) for generating a full trajectory; (iii)-(iv) path quality—including path length and minimum distance to obstacles. against baseline reactive planners based on state-of-the-art field-based methods across a variety of randomized scenarios (Subsection IV-A), and to demonstrate GeoPF’s feasibility and reliability in real robotic experiments with a Franka Emika arm (Subsection IV-B).
To ensure fairness, all experiments were run on the same hardware. [^2] AMD Ryzen 7 5800H, 32GB DDR4-3200 under Ubuntu 20.04, gcc/c++ 9.4. The randomization seeds, number of obstacles, and various obstacle positions for each scenario vary to thoroughly test each method’s robustness.
Unless otherwise noted, we fix the same gains for GeoPF across all trials, highlighting its minimal tuning requirement.
IV-A Quantitative Analysis
This section provides a quantitative assessment of the performance of GeoPF and a comparison against baseline state-of-the-art field planners in a simulated environment. [^3] For the simulated scenarios, a collision is considered when the distance between the robot and an obstacle is less than $10^{-8}$.
  • Potential Field (PF): A standard artificial potential field with point- or sphere-based obstacle approximations;
  • Circular Field (CF): A force-based variant exploiting circulatory vector fields along spherical obstacles [35].
We evaluate the planners under the following scenarios with randomized conditions, and different spherization strategies (optimized to improve PF and CF results). For each scenario, we generate $1000$ cases randomly sampling obstacles. All scenes were defined in a $x,y,z{\in}[{-}0.2,0.2]$ m workspace with initial $[0,-1,0]$ and goal positions $[0,-1,0]$.
  • Line obstacles are generated with multiple randomly generated line obstacles, as seen in Fig. 8. We classify scenarios as “easy” (5–10 lines) or “hard” (10–50 lines);
  • Scenarios with multiple planes are also classified as “easy” or “hard, according to the number of randomly generated planes, 2–8 for the former and 10-40 planes for hard scenarios. For plane-based scenarios, we also consider “longer” scenarios (changing the goal position);
  • Maze-Scenario: designed over classic planning task, we compared the methods over different maze conditions.
The results with key performance metrics, [^4] The minimum distance is the smallest value across the trajectory, while the average takes all the robot-obstacle distances through the whole trajectory. for all planners, are shown in Table II and further detailed below. For the maze scenario, we only present an illustrate case in Fig. 9.
Discussion and Results
For Line Obstacle Scenarios, both the CF and PF require approximating each line segment with small spheres, as seen in Fig. 8. While small spheres ($r^{SP}=0.01$) increase accuracy, they inflate the number of obstacles (and thus computational time). Larger spheres $r^{SP}=0.05$ reduce the count but degrade fidelity. [^5] Since PF is not as competitive as CF [4], we only consider $r^{SP}{=}0.01$. The field parameter for GeoPF was set to $k^{Geo}=0.1$ across all tasks. For PF and CF, we primarily used $k^{attr}=1$, although in some cases, a lower value of $k^{attr}=0.1$ led to improved performance. For each task, we performed a limited parameter search to identify optimal values for PF and CF.
The results, in Table. II, show that GeoPF consistently obtains higher success rates with shorter computational times (up to 8×8$\times$8 × faster). This is partly due to the closed-form distance/force computations for line segments. In contrast, CF and PF often suffer from large overheads due to sphere decompositions, especially in cluttered scenarios.
((a)) Comparison of GeoPF and CF Spherization in Easy (Left) and Hard (Right) Scenarios. The CF parameters in the plot are the one achieves $100\%$ success rate in Table. II First Part. Black lines represent line obstacles, while black circles indicate the spherization of the lines. In cluttered environments, spherization forces the trajectory to navigate around all obstacles. GeoPF plans the trajectory through narrow spaces between obstacles and successfully reaches the goal.
((b)) Spherization of Line and Plane
((c)) Trajectory in Plane Scenario
Figure 8: Comparison of GeoPF and CF Spherization in (a) Ramdom Line and (c) Random Plane Scenarios. (b) illustrates the increased scale of objects (spheres) in spherization from lines to planes, resulting in a sharp increase in computational time. (c) While GeoPF gets a feasible path through obstacles, CF and PF are trapped in local minima, leading to a lower success rate compared to GeoPF in Plane Scenarios.
Figure 9: Maze Task. The robot should find a feasible path with avoiding the plane and go through the tunnel.
TABLE II: Quantitative Analysis and comparison between CF, PF, and GeoPF planners. Key metrics: Obs.N. (Number of obstacles), SR (success rate), CT (s) (total computation time), Step (iterations), CT/Step (s), PL (path length), Min and Av. Dist. (distance to obstacles).
Mean / STD M. ($r^{SP},k^{attr}$) Obs.N. SR CT (s) Step CT/Step (ms) PL (m) Min. Dist. (m) Av. Dist. (m)
[…]
Mean / STD M. ($r^{SP},k^{SP}$) Obs.N. SR CT (s) Step CT/Step (ms) PL (m) Min. Dist. (m) Av. Dist. (m)
[…]
TABLE III: GeoPF performance in complex and dynamic scenarios. Average (black) and standard deviation (gray) shown.
Scenario Obs.N. SR CT (s) Min. Dist. (m)
Complex $12$/$2.45$ 93.7% 3.56/1.37 .032.032.032.032/.03.03.03.03
Dynamic (easy) $\frac{15}{2}(\frac{5}{2})$/$1.1(1.05)$ 94.9% 2.47/0.77 .026.026.026.026/.02.02.02.02
Dynamic (hard) $15(7)$/$1.73(1)$ 83.2% 4.78/1.17 .017.017.017.017/.02.02.02.02
For Plane-based Obstacle Scenarios, we conducted a new parameter search to identify optimal values for PF and CF, as the previously used parameters consistently led to failure. In this evaluation, we analyzed different repulsive coefficients for PF and CF, $k^{SP}=1,10,100$. GeoPF uses the same parameters as in the line task, with $k^{attr}=1$, $k^{Geo}=0.1$.
The results indicate that plane spherization generates a significantly large number of spheres, resulting in an extremely high computational time. Indeed, except for simple line-based scenes, CF and PF planners fail to generate a feasible solution within the real-time framework, as seen in the CT/Step (ms) column in Table II.
In contrast, GeoPF drastically lowers the total computational time—up to 100×100$\times$100 × faster in almost all cases—thanks to direct distance-to-plane calculations. It is also able to keep the real-time performance (computational below 1 ms) for all cases,11 while maintaining a higher success rate in more cluttered arrangements.
The classic Maze Task scenario, where the robot must find a feasible path across multiple planes positioned either vertically or horizontally, is considered a standard trap case. As shown in Fig. 9, while CF and PF fail at the centre point in front of the middle plane, GeoPF successfully navigates through the maze. This result reinforces the findings in Table II which show that CF and PF exhibit significantly lower success in complex scenarios.
Complex and Dynamic Scenes
We additionally tested GeoPF in challenging environments composed of multiple geometric primitives—including lines, planes, cubes, and cylinders—all placed or moving in random configurations. In these trials, only GeoPF was evaluated, as our goal was to demonstrate its versatility in cluttered, complex, and dynamic scenarios.
Specifically, each complex composite scene randomly selected 5–10 lines, 2–5 planes, and 2–3 cubes or cylinders distributed around the workspace. Some obstacles were static, while others followed randomized trajectories (e.g., drifting planes or cylinders). In Table II, the notation Obs. N. is represented as $a(b)$, where $a$ denotes the total number of random objects, while (b) the dynamic ones, including all geometric primitives.
Throughout these tests, GeoPF maintained high success rates without requiring special tuning or re-initialization. The same $k$ value parameters were used. These results confirm the flexibility and easiness to use, as well as the reliability of GeoPF which preserves its real-time efficiency in more complex scenes with a stable and reliable behaviour.
Figure 10: Non-stationary Scenario. The orange plane moves upward over time. The trajectory passes through the initial plane position as shown in the dark orange area. However, since the plane moves upward along the trajectory evolution, it is no longer an obstacle between robot and goal towards the end. Consequently, the robot follows a (straight) line trajectory in last period of the task.
IV-B Real Robot Experiment
To validate the practical viability of GeoPF, we deployed it on a Franka Emika 7-DoF robotic arm. End-effector velocities were generated by mapping the resultant GeoPF forces into joint-space commands using a standard Jacobian pseudoinverse. Obstacles were tracked via a motion-capture system (Nokov Mars) and fitted to corresponding geometric primitives.
GeoPF was evaluated across three representative tasks. First (Fig. 11(a)), a tabletop scenario, where the robot moved from one corner of the table to another while avoiding randomly placed objects. By leveraging the geometric shapes, GeoPF produced smooth, collision-free trajectories under rigid constraints (fixed z-axis), well-suited for control-based execution. The second task (Fig. 11(c)) involved a 3D setting, navigating around a shelf decomposed into multiple planar elements, along with additional obstacles. GeoPF successfully avoided all obstacles by generating environment-aware trajectories. Lastly, in a dynamic scene (Fig. 1), the robot must pour tea into a moving cup while avoiding randomly moving obstacles. Here, GeoPF updated repulsive forces in real time, consistently maintaining safe distances without parameter adjustment. Indeed, across all experiments, GeoPF generated stable, smooth trajectories within real-time cycles without tuning between scenarios.
((a)) 2D Task in Real Robot
((b)) Trajectory for 2D Task
((c)) 3D Task in Real Robot
((d)) Trajectory for 3D Task
Figure 11: Real world experiments. The robot goes from left (low-transparent position) to right (high-opacity) and then returns as seen in trajectories (b),(d). The upper figures show a 2D tabletop task with multiple obstacles, as lines, cubes, and cylinders. The workspace limitation is shown by planes, defining the robot’s boundaries. The bottom figures illustrate a 3D task with four lids representing walls that the robot must avoid during navigation.
V Conclusion
This work introduces a novel reactive planner (GeoPF) that unifies spatial geometric primitives under a temporal potential-field framework for real-time planning. By explicitly leveraging closed-form distance functions for spheres, lines, planes, cubes, and cylinders, GeoPF delivers substantial improvements over existing potential-field methods in computational efficiency, tuning simplicity, and collision avoidance performance. Indeed, quantitative analysis and experimental results demonstrate that GeoPF consistently outperforms existing potential-field-based methods in terms of reliability, ease of deployment, and computational complexity—being the only method suitable for real-time ($1$ kHz) performance in cluttered, unstructured dynamic settings. Future works will involve the integration of prediction modules for obstacle tracking and multi-agent vector fields for effective environment exploration.